# The Impact of Partial Computations on the Red-Blue Pebble Game

### v1

@@ 4.2.1. Zipper gadget
One popular gadget in RBP is the zipper gadget from (Papp and Wattenhofer, 2020; Böhnlein et al., 2025). This consists of two groups of $d$ source nodes, and a long chain where each node has incoming edges from the previous chain node, and one of the two source groups alternatingly.
@@
[…]
Figure 2. Illustration of some frequently used substructures and gadgets: the zipper gadget of (Papp and Wattenhofer, 2020; Böhnlein et al., 2025) with $2\cdot d$ source nodes and an alternating chain (left), a binary tree on $2^{d}$ leaves with all edges pointing towards the root (middle), and a pebble collection gadget from (Papp and Wattenhofer, 2020) with $d$ source nodes and a chain that periodically uses the different sources as an input (right).
@@ 4.2.2. Binary trees
Another structure that often occurs in the context of RBP is a rooted binary (or $k$-ary) tree, i.e. a tree of depth $d$ with every non-leaf node having exactly $2$ (or exactly $k$) distinct in-neighbors. These trees often show up as smaller parts of concrete computational tasks (Saha and Ye, 2024), or as examples to illustrate the properties of different model variants (Böhnlein et al., 2025).
In terms of pebbling, the most interesting behavior for binary trees is observed when $r=3$. Due to their very regular structure, the optimal pebbling strategy is easy to find in these DAGs. Note that we always have to load the $2^{d}$ leaves and save the root node, so the trivial I/O cost is $12^{d}+$ 1. Apart from this, in RBP, for each node $v$ that is not in the bottom two levels (i.e., not a leaf or an out-neighbor of a leaf), we will always need to save (and later reload) one of the two in-neighbors of $v$, in order to free up a red pebble and also compute the other subtree of $v$. This way, the number of non-trivial I/O steps in the optimal RBP pebbling is
@@ 4.2.3. Pebble-collecting gadgets
When proving properties of RBP in a construction, there is often need for a gadget which ensures that we need to place all red pebbles on it simultaneously, or otherwise we incur some I/O cost. One example of such gadgets is the pyramid gadget (Demaine and Liu, 2018; Ranjan et al., 2012), which is similar to a binary tree. A more sophisticated version of the gadget is the one from (Papp and Wattenhofer, 2020), shown in the right in Figure 2. This consist of $d$ source nodes, and a long chain of $\ell$ nodes, with the $i$-th node in the chain having in-edges from the $(i-1)$-th chain node and the $(i\mod d)$-th source node.
This gadget ensures that if we use $d\!+\!2$ red pebbles, then we can pebble it without any I/O cost. However, if there is no point in the pebbling where we have $d\!+\!2$ red pebbles on the gadget simultaneously, then the red pebbles have to be repeatedly reloaded, incurring a cost of at least $\ell/d$.
As a tool for future constructions, we show that the same gadget ensures a similar behavior in PRBP. Note that with $d+2$ red pebbles, the gadget can also be pebbled in PRBP with only the trivial cost.
@@
If a PRBP strategy never has at least $d\!+\!2$ red pebbles simultaneously on this gadget, then its cost is at least $\frac{\ell}{2d}$.
@@ 4.3. On the ratio between $\texttt{OPT}_{RBP}$ and $\texttt{OPT}_{PRBP}$
There exists a DAG construction with $\Delta_{in}=2$ and $\Delta_{out}=3$ such that $\texttt{OPT}_{RBP}=\Theta(n)$ and $\texttt{OPT}_{PRBP}=2$.
@@
The proof uses many copies of the gadget shown before in Figure 1, concatenated in a serial fashion. That is, let us remove the nodes $u_{0}$, $v_{0}$ and the dashed edges from the DAG in the figure. Then consider many copies of the resulting gadget, always merging node $v_{1}$ of the $i$-th gadget with node $u_{1}$ of the $(i\!+\!1)$-th gadget, and node $v_{2}$ of the $i$-th gadget with node $u_{2}$ of the $(i\!+\!1)$-th gadget. In the end, we reinsert the source $u_{0}$ and add two edges towards $u_{1}$ and $u_{2}$ in the first gadget, and reinsert $v_{0}$, and add two edges from $v_{1}$ and $v_{2}$ in the last gadget. Note that the gadget consists of only $8$ nodes, so the resulting DAG contains $\Theta(n)$ copies of the gadget. Assume again that $r=4$.
@@ 4.4. Complexity of $\texttt{OPT}_{RBP}=\texttt{OPT}_{PRBP}$
This is a more technical proof that requires a modified version of a DAG construction from previous works (Papp and Wattenhofer, 2020; Böhnlein et al., 2025). This allows for a reduction from a version of the maximum independent set problem. Recall that given an undirected graph $G_{0}=(V_{0},E_{0})$, an independent set in is a subset of nodes $V^{\prime}\subseteq V_{0}$ that induces no edges, and a maximum independent set is one that maximizes $|V^{\prime}|$ over all possible independent sets in $G_{0}$.
@@
The main ingredient of the reduction is the pebble collection gadget of Proposition 4.6, with a choice of $d=r-2$; we have shown that this works identically in PRBP. Given a maxinset-vertex problem, for each node $u_{0}$ in $G_{0}$, we create two such gadgets, i.e. two groups of $(r-2)$ source nodes, denoted $H_{1}(u_{0})$ and $H_{2}(u_{0})$, and their corresponding chains. Besides this, for some parameter $b$, we merge the $i$-th node of $H_{1}(u_{0})$ and the $i$-th node of $H_{2}(u_{0})$ into a single node for all $i\in[b]$. Finally, for each edge $(u_{1},u_{2})$ in $G_{0}$, we replace a node in $H_{2}(u_{2})$ by a node in the middle of the chain of $H_{1}(u_{1})$, and vice versa, we replace a node in $H_{2}(u_{1})$ by a node in the middle of the chain of $H_{1}(u_{2})$.
With $\ell$ large enough, the gadgets ensure that we need to have all the $r$ red pebbles on them simultaneously in any reasonable solution. In particular, if for every gadget we use all the red pebbles together to compute it, then the total cost in the whole DAG is less than $\frac{\ell}{2(r-2)}$ even in the worst case, whereas if we do not do this for any single gadgets, the cost becomes at least $\frac{\ell}{2(r-2)}$ according to Proposition 4.6. Due to this, any reasonable pebbling strategy comes down to the order of ‘visiting’ the gadgets, i.e. having all red pebbles on them. If $H_{1}(u_{0})$ and $H_{2}(u_{0})$ are visited consecutively, then this saves a cost of $b$ by not having to reload their merged source nodes again. However, the set of groups visited consecutively can only be an independent set in $G_{0}$, due to the dependencies introduced between groups. As such, in both RBP and PRBP, the optimal pebbling visits the gadget-pairs of a largest-possible independent set in $G_{0}$ consecutively. For more details on the main idea of the construction, see  (Papp and Wattenhofer, 2020; Böhnlein et al., 2025).
To extend this reduction idea to our current theorem, consider the input node $v_{0}$ from maxinset-vertex. We further take two sets $Z_{1}$, $Z_{2}$ of $3$ nodes from $H_{1}(v_{0})$ and $H_{2}(v_{0})$, respectively. We add an extra sink node $w$ with edges from all $6$ nodes in $Z_{1}$ and $Z_{2}$. Saving the final value of $w$ incurs an extra trivial cost of $1$, so we disregard this step in the analysis below.
@@ 5. S-partitions and lower bounds
One of the main applications of red-blue pebbling is to develop lower bounds on $\texttt{OPT}_{RBP}$, i.e. the number of required I/O steps for specific computational DAGs from concrete applications. In both the original paper of Hong and Kung (Hong and Kung, 1981) and many follow-up works (Elango et al., 2015; Ranjan et al., 2012), this is done via a special kind of DAG partitioning concept that allows to derive such lower bounds. In this section, we show that the same partitioning concept does not allow us to directly develop bounds that carry over to PRBP.
@@ 5.1. S-partitions in RBP
A subset $D\subseteq V$ is a dominator for $V_{0}$ if every path from a source node to a node in $V_{0}$ contains a node of $D$.
@@
The terminal set of $V_{0}$ is the set of nodes in $V_{0}$ that have no out-neighbors contained in $V_{0}$.
Given these, an $S$-partition is defined as follows.
@@
Hong and Kung show that if we have any valid RBP pebbling in $G$ with $r$ red pebbles, then we can use this to generate a $S$-partition of $G$ with $S=2r$, into $k$ classes such that the I/O-cost of the pebbling strategy is between $r\cdot k$ and $r\cdot(k-1)$. On a high level, this is obtained by cutting up the pebbling strategy into smaller subsequences by splitting after every $r$-th I/O operation; we then create a class $V_{i}$ for each subsequence, and sort each node $v\!\in\!V$ into the class of the first subsequences which places a red pebble on $v$. Intuitively, the $i$-th subsequence starts with at most $r$ red pebbles on $G$, and can load red pebbles on at most $r$ more nodes, so these $2r$ nodes are a dominator for $V_{i}$: all nodes computed in the subsequence indirectly come from these input values. Similarly, the $i$-th subsequence ends with at most $r$ red pebbles on $G$, and saves at most $r$ nodes to slow memory; the terminal set of $V_{i}$ is a subset of these $2r$ nodes. We refer the reader to (Hong and Kung, 1981) for more details.
@@
As such, analyzing the $S$-partitions of concrete DAGs allow us to derive lower bounds on their I/O cost.
@@ 5.2. S-partitions in PRBP
There exists a construction where $\texttt{OPT}_{PRBP}=O(1)$, but we have $\texttt{MIN}_{part}(2r)=\Theta(n)$, i.e., the smallest $S$-partition according to (Hong and Kung, 1981) consists of $\Theta(n)$ classes.
@@
Now consider an $S$-partition for this DAG with $S=6$. The partition cannot consist of a single class, since with $7$ source nodes, there exists no dominator set of size $6$. Let $V^{\prime}$ denote the class that contains the sink node $v$. Note that there is at least one group $H_{i}$ that has none of its nodes in $V^{\prime}$; otherwise, there exists no dominator set of size $6$ for $V^{\prime}$. The nodes in this set $H_{i}$ have no out-neighbors in their corresponding class (since $v\in V^{\prime}$), so all these nodes in $H_{i}$ are contained in the terminal set of their own class. As the terminal sets can have size $6$ at most, this implies that these nodes in $H_{i}$ must span $(|H_{i}|-6)/6$ classes at least. With $|H_{i}|=\Omega(n)$, the $S$-partition then consists of $\Theta(n)$ classes at least. This means that the lower bound obtained from the $S$-partition is also $\Omega(n)$. ∎
@@
Figure 3. Construction for Lemma 5.4, consisting of $7$ source nodes $u_{1},...,u_{7}$, $7$ distinct groups $H_{1},...,H_{7}$ of $\Theta(n)$ nodes each, and a single sink $v$. The node $u_{i}$ always has edges to all the nodes in $H_{i}$, and all the nodes in $H_{i}$ have an edge towards $v$.
@@ 6.1. $S$-edge partitions
We first adapt our auxiliary definitions to an edge-based partitioning. Let $E_{0}\subseteq E$ be a subset of edges in $G$.
@@
A subset of nodes $D\subseteq V$ is an edge-dominator for $E_{0}$ if for every path $\pi$ that starts at a source node and contains an edge in $E_{0}$, it holds that $\pi$ contains a node of $D$.
Note that if $Start(E_{0})$ denotes the starting points in $E_{0}$, i.e. $Start(E_{0})=\{u\in V\,|\,\exists\,v\in V\text{ with }(u,v)\in E_{0}\}$, then $D\subseteq V$ is an edge-dominator for $E_{0}$ if and only if it is a dominator for $Start(E_{0})$.
@@
The edge-terminal set of $E_{0}$ is the set of nodes $V^{\prime}\subseteq V$ with at least one incoming edge in $E_{0}$, but no outgoing edges in $E_{0}$.
@@
An $S$-edge partition of $G$ is a disjoint partitioning $E_{1}$, …, $E_{k}$ of the edges of $E$ of $G$ such that
@@ 6.2. $S$-dominator partitions
For the second approach, we revisit a partitioning concept from (Hong and Kung, 1981).
@@ 6.3.1. Fast Fourier Transform
Among the structured DAGs that describe concrete computations, one of the most frequently studied is the Fast Fourier Transform (FFT). Given some integer $m$ (a power of $2$), the $m$-point FFT graph, also known as the butterfly graph, can be obtained recursively by taking two copies of the $\frac{m}{2}$-point FFT graph, labeling its sink nodes $u_{1}$, …, $u_{m}$, adding a new layer $v_{1}$, …, $v_{m}$, and adding edges from all $u_{i}$ to the nodes $v_{j}$ with $i\equiv j\mod\frac{m}{2}$ (see Figure 4).
@@
The work of (Hong and Kung, 1981) already establishes a tight lower bound of $\Omega\left(\frac{m\log m}{\log r}\right)$ on these graphs via $S$-partitions, and looser bounds are also obtained with spectral methods (Jain and Zaharia, 2020). With our tools above, it is straightforward to show that these bounds also carry over to PRBP.
@@
The proof of the same lower bound for RBP in (Hong and Kung, 1981) actually does not rely on an $S$-partition of the DAG, only on an $S$-dominating partition. In particular, the authors show that

### v2

@@ 4.2.1. Zipper gadget
One popular gadget in RBP is the zipper gadget from (RBpebbling3; mpp). This consists of two groups of $d$ source nodes, and a long chain where each node has incoming edges from the previous chain node, and one of the two source groups alternatingly.
@@
......dd nodesdd nodesdd levels.........dd nodes
Figure 2. Illustration of some frequently used substructures and gadgets: the zipper gadget of (RBpebbling3; mpp) with $2\cdot d$ source nodes and an alternating chain (left), a binary tree on $2^{d}$ leaves with all edges pointing towards the root (middle), and a pebble collection gadget from (RBpebbling3) with $d$ source nodes and a chain that periodically uses the different sources as an input (right).
@@ 4.2.2. Binary trees
Another structure that often occurs in the context of RBP is a rooted binary (or $k$-ary) tree, i.e. a tree of depth $d$ with every non-leaf node having exactly $2$ (or exactly $k$) distinct in-neighbors. These trees often show up as smaller parts of concrete computational tasks (ml1), or as examples to illustrate the properties of different model variants (mpp).
In terms of pebbling, the most interesting behavior for binary trees is observed when $r=3$. Due to their very regular structure, the optimal pebbling strategy is easy to find in these DAGs. Note that we always have to load the $2^{d}$ leaves and save the root node, so the trivial I/O cost is $2^{d}+1$. Apart from this, in RBP, for each node $v$ that is not in the bottom two levels (i.e., not a leaf or an out-neighbor of a leaf), we will always need to save (and later reload) one of the two in-neighbors of $v$, in order to free up a red pebble and also compute the other subtree of $v$. This way, the number of non-trivial I/O steps in the optimal RBP pebbling is
@@ 4.2.3. Pebble-collecting gadgets
When proving properties of RBP in a construction, there is often need for a gadget which ensures that we need to place all red pebbles on it simultaneously, or otherwise we incur some I/O cost. One example of such gadgets is the pyramid gadget (RBpebbling2; ranjan2012upper), which is similar to a binary tree. A more sophisticated version of the gadget is the one from (RBpebbling3), shown in the right side of Figure 2. This gadget consist of $d$ source nodes, and a long chain of $\ell$ nodes, with the $i$-th node in the chain having incoming edges from the previous chain node and the $i$-th source node modulo $d$.
This gadget ensures in RBP that if we use d+2d\!+\!2 red pebbles, then we can pebble it without any I/O (apart from the trivial cost). However, if there is no point in the pebbling where we have d+2d\!+\!2 red pebbles on the gadget simultaneously, then the red pebbles need to be repeatedly reloaded to compute the chain, incurring a cost of at least $\ell/d$.
As a tool for future constructions, we show that the same gadget ensures a similar behavior in PRBP. Note that if we use $d+2$ red pebbles, then the gadget can also be pebbled in PRBP with only the trivial I/O cost. Otherwise, the pebbling again becomes rather costly.
@@
If a PRBP strategy never has at least d+2d\!+\!2 red pebbles simultaneously on this gadget, then the cost of the strategy is at least $\frac{\ell}{2d}$.
@@ 4.3. On the ratio between $\texttt{OPT}_{RBP}$ and $\texttt{OPT}_{PRBP}$
There exists a DAG construction with $\Delta_{in}=2$ and $\Delta_{out}=3$ such that we have $\texttt{OPT}_{RBP}=\Theta(n)$ but $\texttt{OPT}_{PRBP}=2$.
@@
The proof uses many copies of the gadget shown before in Figure 1, concatenated in a serial fashion. That is, let us remove the nodes $u_{0}$, $v_{0}$ and the dashed edges from the DAG in the figure. Then consider many copies of the resulting gadget, always merging node $v_{1}$ of the $i$-th gadget with node $u_{1}$ of the (i+1)(i\!+\!1)-th gadget, and node $v_{2}$ of the $i$-th gadget with node $u_{2}$ of the (i+1)(i\!+\!1)-th gadget. In the end, we reinsert the source $u_{0}$ and add two edges towards $u_{1}$ and $u_{2}$ in the first gadget, and reinsert $v_{0}$, and add two edges from $v_{1}$ and $v_{2}$ in the last gadget. Note that the gadget consists of only $8$ nodes, so the resulting DAG contains $\Theta(n)$ copies of the gadget. Assume again that $r=4$.
@@ 4.4. Complexity of $\texttt{OPT}_{RBP}=\texttt{OPT}_{PRBP}$
This is a more technical proof that requires a modified version of a DAG construction from previous works (RBpebbling3; mpp). This allows for a reduction from a version of the maximum independent set problem. Recall that given an undirected graph $G_{0}=(V_{0},E_{0})$, an independent set is a subset of nodes $V^{\prime}\subseteq V_{0}$ that induces no edges, and a maximum independent set is one that maximizes $|V^{\prime}|$ over all possible independent sets in $G_{0}$.
@@
The main ingredient of the reduction is the pebble collection gadget of Proposition 4.6, with a choice of d=r2d=r-2; we have shown that this works identically in PRBP. Given a maxinset-vertex problem, for each node $u_{0}$ in $G_{0}$, we create two such gadgets, i.e. two groups of (r2)(r-2) source nodes, denoted $H_{1}(u_{0})$ and $H_{2}(u_{0})$, and their corresponding chains. Besides this, for some parameter $b$, we merge the $i$-th node of $H_{1}(u_{0})$ and the $i$-th node of $H_{2}(u_{0})$ into a single node for all $i\in[b]$. Finally, for each edge $(u_{1},u_{2})$ in $G_{0}$, we replace a node in $H_{2}(u_{2})$ by a node in the middle of the chain of $H_{1}(u_{1})$, and vice versa, we replace a node in $H_{2}(u_{1})$ by a node in the middle of the chain of $H_{1}(u_{2})$.
With $\ell$ large enough, the gadgets ensure that we need to have all the $r$ red pebbles on them simultaneously in any reasonable solution. In particular, if for every gadget we use all the red pebbles together to compute it, then the total cost in the whole DAG is less than $\frac{\ell}{2(r-2)}$ even in the worst case, whereas if we do not do this for any single gadgets, the cost becomes at least $\frac{\ell}{2(r-2)}$ according to Proposition 4.6. Due to this, any reasonable pebbling strategy comes down to the order of ‘visiting’ the gadgets, i.e. having all red pebbles on them. If $H_{1}(u_{0})$ and $H_{2}(u_{0})$ are visited consecutively, then this saves a cost of $b$ by not having to reload their merged source nodes again. However, the set of groups visited consecutively can only be an independent set in $G_{0}$, due to the dependencies introduced between groups. As such, in both RBP and PRBP, the optimal pebbling visits the gadget-pairs of a largest-possible independent set in $G_{0}$ consecutively. For more details on the main idea of the construction, see  (RBpebbling3; mpp).
To extend this reduction idea to our current theorem, consider the input node $v_{0}$ from maxinset-vertex. We further take two sets $Z_{1}$, $Z_{2}$ of $3$ nodes from $H_{1}(v_{0})$ and $H_{2}(v_{0})$, respectively. We add an extra sink node $w$ with edges from all $6$ nodes in $Z_{1}$ and $Z_{2}$. Saving the final value of $w$ always incurs an extra trivial cost of $1$; we disregard this step in the analysis below.
@@ 5. S-partitions and lower bounds
One of the main applications of red-blue pebbling is to develop lower bounds on $\texttt{OPT}_{RBP}$, i.e. the number of required I/O steps for specific computational DAGs from concrete applications. In both the original paper of Hong and Kung (RBpebbling1) and many follow-up works (elango2015characterizing; ranjan2012upper), this is done via a special kind of DAG partitioning concept that allows to derive such lower bounds. In this section, we show that the same partitioning concept does not allow us to directly develop bounds that carry over to PRBP.
@@ 5.1. S-partitions in RBP
A subset of nodes $D\subseteq V$ is a dominator for $V_{0}$ if for every directed path $\pi$ that starts at a source node and ends in a node of $V_{0}$, it holds that $\pi$ contains a node of $D$.
@@
The terminal set of $V_{0}$ is the set of nodes $v\in V_{0}$ that satisfy the following property: none of the out-neighbors of $v$ are contained in $V_{0}$.
We note the sets in Definition 5.2 are instead called ‘minimum sets’ in the work of (RBpebbling1); we changed it to terminal sets since we found this original name somewhat misleading.
Given the concepts of dominators and terminal sets, an $S$-partition is defined as follows.
@@
Hong and Kung show that if we have a valid RBP pebbling in $G$ with $r$ red pebbles, then we can use this to generate an $S$-partition of $G$ with $S=2r$ into $k$ classes such that the I/O-cost of the pebbling strategy is between $r\cdot k$ and $r\cdot(k-1)$. On a high level, this is obtained by cutting up the pebbling strategy into smaller subsequences by splitting after every $r$-th I/O operation; we then create a class $V_{i}$ for each subsequence, and sort each node $v\!\in\!V$ into the class of the first subsequences which places a red pebble on $v$. Intuitively, the $i$-th subsequence starts with at most $r$ red pebbles on $G$, and can load red pebbles on at most $r$ more nodes, so these $2r$ nodes are a dominator for $V_{i}$: all nodes computed in the subsequence indirectly come from these input values. Similarly, the $i$-th subsequence ends with at most $r$ red pebbles on $G$, and saves at most $r$ nodes to slow memory; the terminal set of $V_{i}$ is a subset of these $2r$ nodes. We refer the reader to (RBpebbling1) for more details.
@@
As such, analyzing the $S$-partitions of concrete DAGs allows us to derive lower bounds on their I/O cost.
@@ 5.2. S-partitions in PRBP
There exists a construction where $\texttt{OPT}_{PRBP}=O(1)$, but we have $\texttt{MIN}_{part}(2r)=\Theta(n)$, i.e., the smallest $S$-partition according to (RBpebbling1) consists of $\Theta(n)$ classes.
@@
Now consider an $S$-partition for this DAG with $S=6$. The partition cannot consist of a single class, since with $7$ source nodes, there exists no dominator set of size $6$. Let $V^{\prime}$ denote the class that contains the sink node $v$. Note that there is at least one group $H_{i}$ that has none of its nodes in $V^{\prime}$; otherwise, there exists no dominator set of size $6$ for $V^{\prime}$. The nodes in this set $H_{i}$ have no out-neighbors in their corresponding class (since $v\in V^{\prime}$), so all these nodes in $H_{i}$ are contained in the terminal set of their own class. As the terminal sets can have size $6$ at most, this implies that these nodes in $H_{i}$ must span $|H_{i}|/6$ classes at least. With $|H_{i}|=\Omega(n)$, the $S$-partition then consists of $\Theta(n)$ classes at least. This means that the lower bound obtained from the $S$-partition is also $\Omega(n)$. ∎
@@
Figure 3. Construction for Lemma 5.4, consisting of $7$ source nodes $u_{1},...,u_{7}$, 77 distinct groups $H_{1},...,H_{7}$ of $\Theta(n)$ nodes each, and a single sink $v$. The node $u_{i}$ always has edges to all the nodes in $H_{i}$, and all the nodes in $H_{i}$ have an edge towards $v$.
@@ 6.1. $S$-edge partitions
We first adapt our auxiliary definitions to an edge-based partitioning. For this, let $E_{0}\subseteq E$ be a subset of edges in $G$.
@@
A subset of nodes $D\subseteq V$ is an edge-dominator for $E_{0}$ if for every directed path $\pi$ that starts at a source node and contains an edge in $E_{0}$, it holds that $\pi$ contains a node of $D$.
We note that if $Start(E_{0})$ denotes the starting points in $E_{0}$, i.e. $Start(E_{0})=\{u\in V\,|\,\exists\,v\in V\text{ with }(u,v)\in E_{0}\}$, then $D\subseteq V$ is an edge-dominator for $E_{0}$ if and only if it is a dominator for $Start(E_{0})$.
@@
The edge-terminal set of $E_{0}$ is the set of nodes $v\in V$ that satisfy the following property: at least one incoming edge of $v$ is contained in $E_{0}$, but no outgoing edge of $v$ is in $E_{0}$.
@@
An $S$-edge partition of $G$ is a disjoint partitioning $E_{1}$, …, $E_{k}$ of the edges $E$ of $G$ such that
@@ 6.2. $S$-dominator partitions
For the second approach, we revisit a partitioning concept from (RBpebbling1).
@@ 6.3.1. Fast Fourier Transform
Among the structured DAGs that describe concrete computations, one of the most frequently studied is the Fast Fourier Transform (FFT). Consider some integer $m$ that is a power of $2$. The $m$-point FFT graph, also known as the butterfly graph, can be obtained recursively by taking two copies of the $\frac{m}{2}$-point FFT graph, labeling its sink nodes $u_{1}$, …, $u_{m}$, then adding a new layer $v_{1}$, …, $v_{m}$, and adding edges from all $u_{i}$ to the nodes $v_{j}$ with $i\equiv j\mod\frac{m}{2}$. See Figure 4 for an illustration for $m=8$.
@@
The work of (RBpebbling1) already establishes a tight lower bound of $\Omega\left(\frac{m\log m}{\log r}\right)$ on these graphs via $S$-partitions, and looser bounds are also obtained with spectral methods (jain2020spectral). With our tools above, it is straightforward to show that these bounds also carry over to PRBP.
@@
The proof of the same lower bound for RBP in (RBpebbling1) actually does not rely on an $S$-partition of the DAG, only on an $S$-dominating partition. In particular, the authors show that